# k* Distribution: Evaluating the Latent Space of Deep Neural Networks using Local Neighborhood Analysis

### v1

@@ IV-G Effect of Different Distance Metrics on k* Distribution
TABLE XX: Performance using Different Distance Metrics across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$)  (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$)  (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$)  (Clustered) Average of 1,000 ImageNet1k Classes
@@
TABLE XXI: Performance using different tasks and architectures across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and Entire Dataset. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average on Entire Dataset
@@
To evaluate the sensitivity of the nearest neighbor method to different distance metrics and understand their impact on k* distribution and values, we compute the k* distribution using Euclidean ($l_{2}$ norm), City Block ($l_{1}$ norm), Max Norm ($l_{\infty}$ norm), and Cosine Distances. The results from Table XIX reveal the responsiveness of the k* distribution in terms of metrics such as $\mu_{k^{*}}$, $\sigma_{k^{*}}$, and $\gamma_{k^{*}}$. Despite variations, there is substantial agreement on the sample distribution classification across different distance metrics. Having said that, this trend shifts when collectively assessing performance, as shown in Table XX. Here, we observe minimal sensitivity in metric values like $\mu_{k^{*}}$ and $\gamma_{k^{*}}$ to the choice of distance metrics. In other words, the selection of distance metrics does impact the number of classes classified into multiple patterns.
@@ IV-H Going Beyond Image Classification
Exploring beyond the domain of computer vision, we choose to apply k* distribution on models trained for tasks beyond image classification. Specifically, we evaluate models trained for classifying intent from MASSIVE dataset [46] in the domain of natural language processing, and models trained for keyword spotting from the Speech Commands dataset [47] in the domain of speech processing as shown in Table XXI. We note the wider applicability of k* distribution to multiple domains and observe a similar trend of higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$ for better accuracy. Interestingly, we observe that the models trained for intent classification and keyword spotting were less fractured compared to image classification, suggesting that the feature and label association by neural networks for such tasks can be further improved.
@@ V Conclusion
In this article, we introduce the k* distribution, a methodology grounded in the local neighborhood, to assess the distribution of samples in the learned latent space of neural networks. Our experimental findings indicate that the distribution can be primarily categorized into three distinct patterns: Pattern A (🟊normal-$\bigstar$) representing Fractured distribution of samples in latent space: Identified by a positively skewed k* distribution. Pattern B (normal-$\clubsuit$) representing Overlapped distribution of samples in latent space: Characterized by a nearly uniform k* distribution. Pattern C (normal-$\spadesuit$) representing Clustered distribution of samples in latent space: Indicated by a negatively skewed k* distribution. Using k* distribution, we analyzed the learned latent space of different models, with different training data distributions, and of different layers. Further, the learned latent space was tested against various input transformations to understand how the input transformations affect the learned latent space and are associated with the training distribution. We hope this methodology and analysis will help understand the neural networks’ latent spaces and improve the distributions of samples in latent space.

### v2

@@ IV-F Analysis of Latent Space of Different Input Transformations
Adversarial Perturbation:
[…]
The results of adversarial samples on robust and non-robust models are presented in Tables XVII and XVIII. Like other transformations, adversarial perturbations cause also fragmentation in the distribution of samples in the latent space. Although models trained with adversarial techniques exhibit increased robustness and improved transferability as demonstrated by Kotyan et al. [71], they are not completely robust. This is indicated by a gradual rise in the fragmentation with stronger perturbations, making a evident trade-off between robustness and accuracy as highlighted by [70, 72].
In summary, these findings highlight the neural network’s sensitivity to variations in input samples, showcasing that a model perceives transformed samples as substantially different from their original counterparts. The results across various input transformations, including adversarial perturbations, indicate that current models struggle to cluster transformed samples together, interpreting them as distinct from the original sample and each other. Observations also indicate that white-box attacks such as PGD exploit this struggle by further fragmenting the distribution of samples into smaller fractures.
IV-G Effect of Different Distance Metrics on k* Distribution
To evaluate the sensitivity of the nearest neighbor method to different distance metrics and understand their impact on k* distribution and values, we compute the k* distribution using Euclidean ($l_{2}$ norm), City Block ($l_{1}$ norm), Max Norm ($l_{\infty}$ norm), and Cosine Distances. The results from Table XIX reveal the responsiveness of the k* distribution in terms of metrics such as $\mu_{k^{*}}$, $\sigma_{k^{*}}$, and $\gamma_{k^{*}}$. Despite variations, substantial agreement exists on the sample distribution classification across different distance metrics. Having said that, this trend shifts when collectively assessing performance, as shown in Table XX. Here, we observe minimal sensitivity in metric values like $\mu_{k^{*}}$ and $\gamma_{k^{*}}$ to choose distance metrics. In other words, the selection of distance metrics does impact the number of classes classified into multiple patterns.
TABLE XX: Performance using Different Distance Metrics across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$)  (Fractured) Average of Classes with Pattern B ($\clubsuit$)  (Overlapped) Average of Classes with Pattern C ($\spadesuit$)  (Clustered) Average of 1,000 ImageNet1k Classes
@@ IV-G Effect of Different Distance Metrics on k* Distribution
TABLE XXI: Performance using different tasks and architectures across varied visual patterns
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average on Entire Dataset
@@ IV-H Going Beyond Image Classification
Exploring beyond the domain of computer vision, we choose to apply the k* distribution to models trained for tasks other than image classification. Specifically, we evaluate models trained for classifying intent from MASSIVE dataset [64] in the domain of natural language processing, and models trained for keyword spotting from the Speech Commands dataset [65] in the domain of speech processing as shown in Table XXI. We note the wider applicability of k* distribution to multiple domains and observe a similar trend of higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$ for better accuracy. Interestingly, we observe that the models trained for intent classification and keyword spotting were less fractured than image classification, suggesting that neural networks’ feature and label association for such tasks can be further improved.
@@ V Conclusion
In this article, we introduce the k* distribution, a methodology grounded in the local neighborhood, to assess the distribution of samples in the learned latent space of neural networks. Our experimental findings indicate that the distribution can be primarily categorized into three distinct patterns: Pattern A (🟊$\bigstar$) representing Fractured distribution of samples in latent space: Identified by a positively skewed k* distribution. Pattern B ($\clubsuit$) representing Overlapped distribution of samples in latent space: Characterized by a nearly uniform k* distribution. Pattern C ($\spadesuit$) representing Clustered distribution of samples in latent space: Indicated by a negatively skewed k* distribution. Using k* distribution, we analyzed the learned latent space of different models, with different training data distributions, and of different layers. Further, the learned latent space was tested against various input transformations to understand how the input transformations affect the learned latent space and are associated with the training distribution. We hope this methodology and analysis will help understand the neural networks’ latent spaces and improve the distributions of samples in latent space.